# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ B.5 Computational Complexity
Table 6: Quantitative results on sampling efficiency of continuous and discrete samplers. PDM/LDM denote pixel-/latent-space continuous diffusion models, respectively; VQ-Diffusion and MMaDA are discrete diffusion models. Rows shaded gray report runtimes on a single NVIDIA A6000 GPU copied from G2D2 (Murata et al., 2024); rows shaded orange are measured by us on a single NVIDIA A100 GPU.
@@
Analysis. Table 6 highlights the sampling efficiency of our discrete posterior sampler (APS) relative to both continuous-diffusion baselines and the prior discrete sampler G2D2.
@@
Against prior discrete diffusion. Under the same budget, APS matches or improves G2D2’s runtime while yielding better reconstructions: at $256\times 256$ with VQ-Diffusion both methods use 100 steps, but APS yields higher quality (see Table 5); with the stronger MMaDA prior at $512\times 512$, APS achieves consistently better LPIPS/PSNR as discussed in §4. Moreover, our $\times1024$ configuration (APS-L) demonstrates better test-time scaling behavior compared continuous diffusion: we keep 15 steps and still obtain substantial quality gains at reasonable cost (484 s).
@@
B.6 Compared Baselines
We compare our method against state-of-the-art posterior samplers using pixel-/latent-space continuous and discrete diffusion models. Each baseline is evaluated under the same data as ours. We follow the experimental setup from G2D2 and reuse the baseline implementations to ensure a fair comparison. To address the resolution mismatch between our model and the benchmark datasets, we adopt the protocol from G2D2 (Murata et al., 2024). The benchmark images are first upsampled to match the input resolution of our base model MMaDA (Yang et al., 2025). The forward corruption operator and our posterior sampling method are then applied in this high-resolution space. Finally, the resulting output is downsampled to the original $256\times 256$ resolution for a fair evaluation. A brief description of each baseline and links to available source code are provided below:
@@ B.6 Compared Baselines
  • G2D2 (Murata et al., 2024): A discrete diffusion posterior sampler that uses a star-shaped noising process and Gumbel-Softmax continuous relaxation to enable gradient guidance in discrete space. We use the source code with the exact implementation and hyperparameters provided in the original paper. Specifically, we use 100 reverse diffusion steps and 30 optimization steps per reverse step. The learning rate and the coefficient for the KL-divergence loss are scheduled logarithmically, as proposed. Source: https://github.com/sony/g2d2
@@ B.7 Benchmarks & Metrics
B.8 Additional Results
B.8.1 General Inverse Problems
@@
Figure 7: Additional qualitative results for super resolution ($4\times$) on ImageNet. Compared to continuous baselines (DPS, DDRM, PSLD, ReSample) and the discrete baseline G2D2, APS produces sharper details and more faithful reconstructions across diverse examples. For instance, in the second row, the leopard’s eyes are reconstructed with finer details.
@@
B.8.2 Text-Guided Large Block Inpainting
@@
Figure 5 shows that our method generates realistic clothing textures, with better alignment to the reference prompts and fewer artifacts. APS leverages discrete diffusion’s ability to directly reweight categorical distributions under posterior guidance, that helps generate visually appealing and semantically accurate completions.
B.8.3 Stylization
@@
Figure 13 provides a qualitative comparison of our full method (MMaDA + APS) against the base model (MMaDA) and state-of-the-art continuous diffusion approaches. This experiment is designed to demonstrate the novel capability of discrete diffusion models for challenging nonlinear style transfer, not solely to outperform continuous alternatives. We observe that competing methods struggle to balance style fidelity with content alignment. Training-free methods like StyleAligned (Hertz et al., 2023) and InstantStyle (Wang et al., 2024) often drift towards generic textures. Conversely, the training-based StyleDrop (Sohn et al., 2023) tends to overfit to superficial color patterns, which compromises semantic coherence with the text prompt. Our base model, MMaDA (Yang et al., 2025), maintains reasonable content fidelity but fails to transfer fine-grained style attributes, such as material textures or stroke-level details. In contrast, our full method (MMaDA + APS) consistently produces outputs that preserve the reference style while maintaining strong semantic alignment. For instance, our result for the “letter” prompt retains the intricate, flowing smoke design, while the “milkshake” example accurately captures the specified retro diner aesthetic. These results highlight the effectiveness of APS in discrete diffusion models for complex style transfer tasks.
@@
B.8.4 Additional Quantitative Results
We perform a larger-scale evaluation for $4\times$ super resolution on FFHQ, extending our analysis to 1000 samples. Our results are compared against the numbers reported for the same task in G2D2 (Murata et al., 2024). Importantly, our observation in the main draft extends to the larger-scale setting and our APS algorithm consistently outperforms G2D2 in all metrics.
@@ B.8.4 Additional Quantitative Results
Discrete G2D2 w/ star-shaped noise process 0.265 27.29 0.763
@@ B.9 Limitations
Failure Cases.
Figure 14 illustrates failure cases of our approach in stylization. We observe that APS sometimes produces over-smoothed outputs when the reference style is out-of-distribution, or introduces artifacts when the measurement operator is poorly aligned with the pretrained backbone. These examples highlight opportunities for improving robustness and generalization.

### v2

@@ C.5.1 Tokenwise Expectation to Differentiable Surrogate
Table 8: Quantitative results on sampling efficiency of continuous and discrete samplers. PDM/LDM denote pixel-/latent-space continuous diffusion models, respectively; VQ-Diffusion and MMaDA are discrete diffusion models. Rows shaded gray report runtimes on a single NVIDIA A6000 GPU copied from G2D2 (Murata et al., 2024); rows shaded orange are measured by us on a single NVIDIA A100 GPU. APS matches or improves the efficiency of prior discrete samplers while achieving significantly better results (§4). With only 15 denoising steps, APS achieves comparable or better reconstruction quality at up to $6\times$ lower runtime than latent diffusion baselines that require 1000 steps.
@@
C.5.2 Runtime Analysis of Anchored Posterior Sampling
Table 8 highlights the sampling efficiency of our anchored posterior sampler (APS) relative to both continuous-diffusion baselines and the prior discrete diffusion sampler G2D2.
@@ C.5.2 Runtime Analysis of Anchored Posterior Sampling
Against prior discrete diffusion (G2D2). Under the same budget, APS matches or improves G2D2’s runtime while yielding better reconstructions: at $256\times 256$ with VQ-Diffusion both methods use 100 steps, but APS yields higher quality as given in Table 7. Moreover, our $\times1024$ configuration (APS-L) demonstrates better test-time scaling behavior compared continuous diffusion: we keep 15 steps and still obtain substantial quality gains at reasonable cost (484 s).
@@
Table 9: Noise robustness analysis for $\times256$ Super-Resolution (4x) on FFHQ. APS consistently outperforms G2D2 across varying noise levels ($\sigma\in\{0.01,0.02,0.03,0.04,0.05\}$) in PSNR, SSIM, and LPIPS. The performance gain of APS remains consistent across different noise levels.
Measurement Noise ($\sigma$) Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0.01 G2D2 25.98 0.739 0.328
APS 27.52 0.784 0.268
0.02 G2D2 25.90 0.738 0.333
APS 27.42 0.781 0.273
0.03 G2D2 25.88 0.733 0.338
APS 27.31 0.774 0.281
0.04 G2D2 25.64 0.726 0.341
APS 27.04 0.764 0.290
0.05 G2D2 25.46 0.717 0.350
APS 26.80 0.759 0.310
C.6 Analysis of Robustness to Measurement Noise
Table 9 evaluates the noise robustness of APS against G2D2 under Gaussian noise with varying standard deviations. APS consistently improves PSNR and SSIM while reducing LPIPS across all noise levels, indicating superior visual fidelity and structural consistency. The performance gap remains stable as $\sigma$ increases, showing that APS maintains high reconstruction quality even under strong measurement corruption. These results, along with experiments spanning different tokenizers (LFQ, VQ-VAE), datasets (FFHQ, ImageNet, StyleAligned), and tasks (inpainting, deblurring, super-resolution), demonstrate the robustness and versatility of our approach.
C.7 Compared Baselines
We compare our method against state-of-the-art posterior samplers using pixel-/latent-space continuous and discrete diffusion models. Each baseline is evaluated under the same data as ours. We follow the experimental setup from G2D2 and reuse the baseline implementations to ensure a fair comparison. To address the resolution mismatch between our model and the benchmark datasets, we adopt the protocol from G2D2 (Murata et al., 2024). The benchmark images are first upsampled to match the input resolution of our base model MMaDA (Yang et al., 2025b). The forward corruption operator and our posterior sampling method are then applied in this high-resolution space. Finally, the resulting output is downsampled to the original $256\times 256$ resolution for a fair evaluation. A brief description of each baseline and links to available source code are provided below:
@@ C.7 Compared Baselines
  • G2D2 (Murata et al., 2024): A discrete diffusion posterior sampler that uses a star-shaped noising process and Gumbel-Softmax continuous relaxation to enable gradient guidance in discrete space. We follow the exact implementation and hyperparameters provided in the original paper. Specifically, we use 100 reverse diffusion steps and 30 optimization steps per reverse step. The learning rate and the coefficient for the KL-divergence loss are scheduled logarithmically, as proposed. Source: https://github.com/sony/g2d2
@@ C.8 Benchmarks & Metrics
Question Answering (Qwen2.5-3B(Yang et al., 2025a): For question answering, we use the Qwen2.5-3B model to conduct a targeted evaluation on open-ended generation without additional training. We curate a benchmark of 20 prompts (listed in Table 10) designed to test various capabilities, including creative writing, list creation, and constrained generation (e.g., specific sentence counts). Evaluation is performed using a dual strategy: Qwen3-0.6B serves as a reward model to guide diffusion inference, while Qwen3-8B acts as an independent oracle judge to assess instruction following capability and semantic coherence.
Table 10: Question Answering Benchmark Prompts. The evaluation set consists of 20 diverse instructions, including creative writing, list generation, and constrained formatting tasks created using Qwen2.5-3B.
ID Prompt ID Prompt
1 Write a short paragraph describing your favorite food. 11 Enumerate three favorite hobbies you have incorporated into your weekend routine.
2 Create a list of three things you love most about your weekend. 12 Write one sentence describing what you did last Saturday.
3 Write a two-sentence story starting with ‘Once upon a time, I found a mysterious letter in my mailbox.’ 13 Create a sentence about your day, detailing what you did at home.
4 Provide reasons why you like singing. 14 Write about one day in your life, focusing on what you watched on TV.
5 Write about your pet’s daily routine. 15 Create a list of three things you like about your city.
6 Explain in two sentences why spending time outdoors is important. 16 Describe your favorite book in one sentence.
7 Create a list of three things you did last Saturday. 17 Write a short summary (5 sentences) of a movie you recently watched.
8 Write down your dream for your dream job, describing what you see in it. 18 Describe one thing you did today that made you happy.
9 Create a simple diary entry about your best day at school. 19 Write (4 sentences) why home cooking is important to you.
10 Write a paragraph that explains why studying is beneficial. 20 Create a paragraph (5 sentences) about your favorite day in the park.
C.9 Additional Results
C.9.1 General Inverse Problems
@@
Figure 7: Additional qualitative results for super resolution ($4\times$) on ImageNet. Compared to continuous baselines (DPS, DDRM, PSLD, ReSample) and the discrete baseline G2D2, APS produces sharper details and more faithful reconstructions across diverse examples. For instance, in the second row, the leopard’s eyes are reconstructed with fine detail, noticeably better than the baselines.
@@
C.9.2 Text-Guided Large Block Inpainting
@@
Comparison.
We compare APS against large-scale continuous diffusion baselines: Imagen3 (Baldridge et al., 2024), Flux (Black Forest Labs, 2024), and HDPainter (Manukyan et al., 2025). As illustrated in Figure 12, our method generates realistic clothing textures, with stronger alignment to the reference prompts and fewer artifacts (red boxes highlight failure regions of competing methods). In contrast, Imagen3 and Flux often introduce distorted or inconsistent garment regions, while HD-painter produces less faithful completions with mismatched styles. APS leverages discrete diffusion’s ability to directly reweight categorical distributions under posterior guidance, yielding visually compelling and semantically accurate completions.
[…]
C.9.3 Stylization
@@
Figure 14 provides a qualitative comparison of our full method (MMaDA + APS) against the base model (MMaDA) and state-of-the-art continuous diffusion approaches. This experiment is designed to demonstrate the novel capability of discrete diffusion models for challenging nonlinear style transfer, not solely to outperform continuous alternatives. We observe that competing methods struggle to balance style fidelity with content alignment. Training-free methods like StyleAligned (Hertz et al., 2023) and InstantStyle (Wang et al., 2024) often drift towards generic textures. Conversely, the training-based StyleDrop (Sohn et al., 2023) tends to overfit to superficial color patterns, which compromises semantic coherence with the text prompt. Our base model, MMaDA (Yang et al., 2025b), maintains reasonable content fidelity but fails to transfer fine-grained style attributes, such as material textures or stroke-level details. In contrast, our full method (MMaDA + APS) consistently produces outputs that preserve the reference style while maintaining strong semantic alignment. For instance, our result for the “letter” prompt retains the intricate, flowing smoke design, while the “milkshake” example accurately captures the specified retro diner aesthetic. These results highlight the effectiveness of anchored posterior sampling in discrete diffusion models for complex style transfer tasks.
@@
C.9.4 Additional Quantitative Results
We perform a larger-scale evaluation for $4\times$ super resolution on FFHQ, extending our analysis to 1000 samples. Our results are compared against the numbers reported for the same task in G2D2 (Murata et al., 2024). Importantly, our observation in the main draft extends to the larger-scale setting and our APS algorithm consistently outperforms G2D2 in all metrics, as given in Table 11.
@@ C.9.4 Additional Quantitative Results
Discrete G2D2 0.265 27.29 0.763
@@ C.10 Limitations
  4. 4.
    Approximate posterior sampling. APS optimizes a surrogate variational objective and does not guarantee asymptotically exact posterior samples. While this approximation is essential for tractability in high-dimensional discrete spaces, it may introduce bias relative to exact discrete MCMC methods, which remain computationally infeasible at scale.
Failure Cases. Figure 15 illustrates failure cases of our approach in stylization. We observe that APS sometimes produces over-smoothed outputs when the reference style is out-of-distribution, or introduces artifacts when the measurement operator is poorly aligned with the pretrained backbone. These examples highlight opportunities for improving robustness and generalization.
@@
C.11 Reproducibility Statement
Our experiments are built upon the publicly available MMaDA codebase (Yang et al., 2025b). All modifications and implementation details are described in Appendix C.1, which includes the pseudocode in Algorithm 1 and the specific parameters used for every experiment. Furthermore, Appendix C.2 provides comprehensive ablation studies and hyperparameter sweeps (Table 5). The experiments utilize the widely-used public datasets FFHQ (Karras et al., 2019) and ImageNet (Deng et al., 2009). The combination of a public codebase and datasets, along with our detailed Algorithm 1 and parameters (§C.2), should ensure that our results are readily reproducible.